OpenStudy (anonymous):

-10-x(15+3x^2) for =-1

OpenStudy (anonymous):

8

OpenStudy (anonymous):

For x=-1 right

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how did you get 8

OpenStudy (anonymous):

-10 - (-1)(15+3(-1)^2) -10 + (15+3) =8 because -1^2=1 - (-1)=1

OpenStudy (anonymous):

(-1)^2 sorry the brackets are important